Deborah has 33 g of H2 in a balloon. The pressure inside the balloon is 1.87 atm and the temperature inside the ballon is 25 Celsius. What's is the volume of the balloon? (Celclius should be converted to Kelvin)
1
Expert's answer
2017-11-18T12:38:07-0500
We use the ideal gas law PV = nRT, where:
• P is the pressure of the gas,
• V is the volume of the gas,
• n is the amount of substance of gas (in moles),
• R is the universal gas constant,
• T is is the absolute temperature of the gas (in kelvins).
P = 1,87 atm = 1,87 × 101325 Pa = 189477,75 Pa;
n = m/M; M = 2 g/mol;
n=33/2=11,5 mol;
R = 8,314 m3⋅Pa⋅K-1⋅mol-1;
T = 273,15 + t = 273,15 + 25 = 298,15 K
V = nRT/P = 11,5⋅8,314⋅298,15/189477,75 = 0,15 m-3 = 150 L.

Answer:
V = 150 L.